OpenStudy (anonymous):

On a particular Saturday, 60% of the visitors to an art museum were students. Iff 144 students visited the museum, how many total visitors did the art museum had Saturday?????????? Will be fan and rewarded a medal if solved.

OpenStudy (anonymous):

144 divide by 6 multiply by 10 in that order

OpenStudy (anonymous):

240 is the answer

OpenStudy (anonymous):

Basically, the equation you are trying to solve is: 144 is 60% of what number. \[\frac{ 144 }{ x } \times \frac{ 60 }{ 100 }\] So, the formula I used is: \[\frac{ is }{ of } x \frac{ % }{ 100 }\] Is stands for how much of that number you know, and of stnads for the toal number.

OpenStudy (anonymous):

It is supposed to be % over 100 for the blank formula

OpenStudy (anonymous):

if u want to put it the scientific way that would be the answer ty mangorox

OpenStudy (anonymous):

can u give me another question similar to this

OpenStudy (anonymous):

After that you cross multiply: \[144 \times 100 = 14,400\] \[60 \times x = 60x\] Then, you divide by 60 on both sides: \[14,400 \div 60 = 240\] \[60x \div 60 = 1\] \[ \frac{ 240 }{ 1 } = 240\] \[x = 240\]

OpenStudy (anonymous):

can u give me another question similar to this

OpenStudy (anonymous):

sorry?

OpenStudy (anonymous):

can u give me another question similar to this

OpenStudy (anonymous):

Ok, try this one: On a particular Monday, 75% of the visitors to an art museum were students. If 90 students visited the museum, how many total visitors did the art museum had Monday?

OpenStudy (anonymous):

So, do you know how to set up the equation to solve this?

OpenStudy (sleepyjess):

If you have 2 parts of the equation, then write it in the form that makes sense for you to begin with. You know you have 144 students and that they make up 60% of the total. So 144 = 60%(Total) Get Total by itself by dividing both sides by 60% (or .60) 144/.60 = Total 240 = Total